Use the discriminant to determine the number and type of solutions of the equations.Your anwser should be one real root, two real roots or two complex roots. Circle the discriminant value.
x^2+8x-6=0
4x^2-4x+1=0
Solve
x^4-9x^2-112=0
x^2/3-7x^1/3+10=0

Answer by stanbon(75887)

Use the discriminant to determine the number and type of solutions of the equations.Your answer should be one real root, two real roots or two complex roots.
----------------------------
Circle the discriminant value.
x^2+8x-6=0
b^2-4ac = 8^2-4*1*-6 >0 so there are 2 real roots
----------------------
4x^2-4x+1=0
b^2-4ac = 4^2-4*4*1 = 0 so there are 2 equal real roots
----------------------
Solve
Use the quadratic formula.
x^4-9x^2-112=0
x^2/3-7x^1/3+10=0
-------------
Cheers,
Stan H.